OpenStudy (anonymous):

Cashews cost $4 per pound and peanuts cost $1 per pound. How many pounds of peanuts should be used with the cashews to get a mixture of 24 pounds that will cost $2 per pound? 16 9 12.5 6

OpenStudy (anonymous):

16

OpenStudy (anonymous):

Assume x pounds of peanuts for x dollars so 24-x pounds of cashews for 4(24-x) then x + 4(24-x) = 48 (cost of mix) 3x = 48 -> x = 16

OpenStudy (amistre64):

while substitution is the usual route, elimination can be simpler to process for some. c + p = 24 ; *-4 4c +1p = 2(24) -4c - 4p = -96 4c +1p = 48 -------------- -3p = -48 ; solve for p p = 48/3 = 16